40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 10 02:15:36 2014, Jenne, Update, LSC, QPD dark noise checkout 
    Reply  Fri Jul 11 20:05:34 2014, Jenne, Update, LSC, QPD dark noise checkout TransQPDY_darkNoise_vs_TransimpedanceGain_9July2014.pdfTransQPDY_darkNoise_vs_WhiteningGain_9July2014.pdfsqrtTrans_sensitivity.pngsqrtTrans_Noise_vs_CARMoffset.png
       Reply  Tue Jul 15 21:43:28 2014, Jenne, Update, LSC, QPDs back 
Message ID: 10174     Entry time: Thu Jul 10 02:15:36 2014     Reply to this: 10187
Author: Jenne 
Type: Update 
Category: LSC 
Subject: QPD dark noise checkout 

I have put beam dumps in front of both of the end transmission QPDs so that I could measure the dark noise.  They are still there.

I checked that the Yend QPD sliders and switches were doing things as I expected.  I couldn't do the Xend since c1auxex is still lost (elog 10165).  I'll post plots and actual information on this checkout, as well as my calculation of what this dark noise means in terms of meters for CARM when we're using 1/sqrt(trans) signals tomorrow morning. 

ELOG V3.1.3-